Please provide step by step answer for this JAVA program. I need it in details. (The correct answer is "10")

int[] numbers = {1,4,5,0,7,8};
int k = 0;
int sum = 0;
while(numbers[k] != 0){
     sum += numbers[k];
     k++;
}

What is the output of the above code?

  

25

   

10

   

0
Answer #1
numbers[0] = 1
sum = sum + 1 = 1

numbers[1] = 4
sum = sum + 4 = 1 + 4 = 5

numbers[2] = 5
sum = sum + 5 = 5 + 5 = 10

numbers[3] = 0
So, the condition of the while loop breaks.

So, the final value of sum is 10